# Data-driven Conditional Instrumental Variables for Debiasing Recommender Systems

### v1

@@ Abstract
In recommender systems, latent variables can cause user-item interaction data to deviate from true user preferences. This biased data is then used to train recommendation models, further amplifying the bias and ultimately compromising both recommendation accuracy and user satisfaction. Instrumental Variable (IV) methods are effective tools for addressing the confounding bias introduced by latent variables; however, identifying a valid IV is often challenging. To overcome this issue, we propose a novel data-driven conditional IV (CIV) debiasing method for recommender systems, called CIV4Rec. CIV4Rec automatically generates valid CIVs and their corresponding conditioning sets directly from interaction data, significantly reducing the complexity of IV selection while effectively mitigating the confounding bias caused by latent variables in recommender systems. Specifically, CIV4Rec leverages a variational autoencoder (VAE) to generate the representations of the CIV and its conditional set from interaction data, followed by the application of least squares to derive causal representations for click prediction. Extensive experiments on two real-world datasets, Movielens-10M and Douban-Movie, demonstrate that our CIV4Rec successfully identifies valid CIVs, effectively reduces bias, and consequently improves recommendation accuracy.
Introduction
@@
With the rapid development of the Internet, the amount of information has exploded, making it increasingly difficult for users to sift through vast amounts of data to find content that aligns with their preferences (Gao et al. 2021). Recommender systems have emerged as a critical solution to this problem by analysing user behaviour data to deliver personalised recommendations, thereby enhancing user engagement and satisfaction (Wang et al. 2020). These systems have become an integral part of many digital platforms, finding extensive applications across various domains, including e-commerce (Shoja and Tabrizi 2019), streaming media (Gomez-Uribe and Hunt 2015), and social networks (Liao et al. 2022), significantly improving information retrieval efficiency and user experience.
@@
Figure 1: Three causal graphs illustrate the assumptions underlying existing work. The variables are defined as follows: $U$: user preferences, $I$: displayed items, $Y$: user feedback or outcome variables, $C_{i}$: confounding factors affecting items, $C$: confounding factors, $Z$: instrumental variables, $W$: treatment variables (i.e., embeddings of user-item pairs). (a) The causal graph represents the traditional recommendation model; (b) the causal graph represents the debiasing method using causality; (c) the causal graph representation of the debiasing method using standard IV.
@@
User behaviour data is crucial for recommender systems in predicting user preferences. Existing models often assume this data is unbiased and accurately reflects user preferences, meaning the data accurately reflects user preferences. Based on this assumption, many methods have been proposed, such as Matrix Factorisation (MF) (Koren, Bell, and Volinsky 2009) and Neural Network-based Collaborative Filtering (NCF) (He et al. 2017). These methods achieve the goal of predicting user preferences by fitting user behaviour data, as depicted in the causal directed acyclic graph (DAG) (Pearl 2009) shown in Figure 1 (a), which illustrates the causal relationships between user preferences $\left(U\right)$ and user feedback $\left(Y\right)$, as well as between displayed items $\left(I\right)$ and $Y$. However, in the real-world, user behaviour is inevitably influenced by various unobservable confounding factors, such as item popularity (e.g., frequent recommendation of an item limits users’ choices) and user psychology (e.g., choosing to watch a movie for socialising) (Chen et al. 2023). These factors introduce biases like popularity and conformity bias, leading models to learn false correlations and reducing their ability to accurately predict preferences.
Causal inference techniques have been utilised to mitigate bias in recommender systems (Wang et al. 2020; Gao et al. 2024). These methods leverage domain knowledge to design causal DAGs that t represent the data generation process, identify specific biases, and develop models to address them. For example, Zhang et al. (Zhang et al. 2021) proposed a novel causal graph (Figure 1 (b)) that analyses how item popularity affects recommendations. Based on this graph, they introduced the Popularity-bias Deconfounding and Adjusting (PDA) training paradigm to mitigate popularity bias. Similarly, Zheng et al. (Zheng et al. 2021) examined the impact of users’ conformity mentality and proposed the Disentangling Interest and Conformity with Causal Embedding (DICE) method to address conformity bias. While these causal graph-based approaches have shown some success, they rely on the assumption that the data generation process strictly follows the proposed causal graph. If real-world conditions deviate significantly from these assumptions, their effectiveness may be limited (Cai et al. 2024). Moreover, the presence of numerous confounding factors in real-world scenarios makes it challenging to satisfy these assumptions.
@@
Latent confounding factors pose a significant challenge in debiasing recommender systems. These latent factors (a.k.a. unobserved or unmeasured variables) influence both the treatment (e.g., the recommendation process) and the outcome, leading to biased recommendations. Instrumental Variables (IVs) are often used to eliminate bias caused by latent variables (Caner and Hansen 2004; Pearl 2009), i.e., $Z$ is an IV as shown in Figure 1 (c). A standard IV must meet three conditions (Pearl 2009): (i) relevance to the treatment variable; (ii) an exclusive impact on the outcome through the treatment; and (iii) no shared confounders with the outcome. Several debiasing methods for recommender systems based on standard IVs have been developed to address latent confounding factors without relying on strict causal graph assumptions (Si et al. 2022, 2023a, 2023b). For instance, IV4Rec (Si et al. 2022) uses self-collected search data as an IV to effectively mitigate latent confounding. However, verifying the last two conditions of a standard IV using data alone is infeasible (Brito and Pearl 2012; Cheng et al. 2023), making it challenging to identify valid IVs.
@@
Figure 2: A causal DAG illustrating our proposed CIV4Rec method for discovering CIVs and their conditional sets from observational data. $W$, $Y$, $X$, and $U_{c}$ are the treatment, outcome, set of measured pretreatment variables, and latent confounders between $W$ and $Y$, respectively. $Z_{t}$ and $Z_{c}$ denote the representations of the CIVs and their conditional sets learned from $X$.
@@
In causal inference, researchers have used conditional IVs (CIVs) (definitions provided in the Appendix) to tackle the limitations of standard IVs (Pearl 2009; Brito and Pearl 2012). CIVs offer more relaxed application conditions than standard IVs. Recently, Cheng et al. (Cheng et al. 2023) developed a CIV method (CIV.VAE) based on the variational autoencoder (VAE) (Kingma and Welling 2013; Schölkopf et al. 2021; Schölkopf 2022) model, which generates CIVs and their conditional sets from data, significantly relaxing the constraints of standard IVs. However, this method is designed for tabular data, and no attempt has been made to adapt it to interactive data in recommender systems.
@@
To address the challenge of using CIV in interactive data, we first propose a causal DAG, as shown in Figure 2, to represent the causal relationships between observed and latent variables in interaction data within recommender systems. Building on this causal DAG, we develop a novel data-driven CIV debiasing method for recommender systems, termed CIV4Rec. Specifically, CIV4Rec uses the embeddings of user-item pairs (including users with the selected items, i.e., positive samples, and users with their pre-selected items, i.e., negative samples) as the treatment variable $W$, the user feedback as the outcome $Y$, and the user interaction data (only positive samples) as the pretreatment variable $X$. We assume that at least one CIV exists within $X$, capturing latent information such as user search behaviours that lead to interactions. CIV4Rec employs a VAE to generate the representations of the CIV and its conditional set from $X$, denoted as $Z_{t}$ and $Z_{c}$, as shown in Figure 2. The contribution of our work is summarised below:
@@
  • We develop a novel data-driven CIV method (CIV4Rec), for learning representations of CIVs and its conditional sets under the proposed causal DAG. To the best of our knowledge, this is the first work to generate representations of CIVs and their conditional sets from interaction data for mitigating bias in recommender systems.
@@
  • Extensive experiments on two real-world datasets demonstrate that CIV4Rec achieves optimal debiasing results and recommendation performance compared to state-of-the-art causal debiasing methods.
Related Work
@@
In this section, we review the recommender methods most closely related to our CIV4Rec, including traditional recommender methods and causal recommender methods.
Traditional Recommender Methods
@@
Traditional recommender methods, primarily based on Collaborative Filtering (CF), typically assume that user behaviour data is unbiased (Koren, Rendle, and Bell 2021). The mainstream approach is model-based CF, which trains a model on user behaviour data to recommend items that align with user preferences. A classic method is MF (Koren, Bell, and Volinsky 2009), which decomposes the user-item rating matrix to predict preferences. However, MF assumes that unselected items are incompatible with user preferences, ignoring cases where users may not have encountered those items. To address this, Rendle et al. (Rendle et al. 2012) proposed Bayesian Personalized Ranking (BPR), which assumes users prefer selected items over unselected ones, enhancing preference inference in MF.
@@
With the rise of deep learning, He et al.(He et al. 2017) introduced NCF, using multi-layer perceptrons (MLPs) to better model non-linear user preferences and improve recommendation accuracy. Researchers also integrated graph structures to capture richer behaviour data. Wang et al.(Wang et al. 2019) developed Neural Graph Collaborative Filtering (NGCF), leveraging Graph Convolutional Networks (GCNs) to enhance recommendations by embedding user-item interactions. Building on NGCF, He et al. (He et al. 2020) proposed LightGCN, a simplified model that improves efficiency and performance. Despite their success, traditional methods often overlook biases (e.g., popularity bias), which can be amplified during training, leading models to overemphasize popular items.
Causal Recommender Methods
@@
To mitigate biases in recommender systems, researchers have increasingly adopted causal inference techniques. Early approaches used the Inverse Propensity Score (IPS) (Wang et al. 2021; Schnabel et al. 2016; Bottou et al. 2013) to reduce bias by assigning an inverse propensity score (e.g., the inverse of item popularity) to user-item interactions during training, balancing the influence of popular and less popular items. However, IPS methods often suffer from high variance and instability. Building on IPS success, researchers have explored causal graph-based methods (Zheng et al. 2021; He et al. 2023; Zhang et al. 2021) that model the generation mechanisms of user behaviour. These methods design specific models to address biases like popularity and conformity bias. Yet, the presence of unobserved confounders in real-world data limits the effectiveness of these causal graph assumptions (Cai et al. 2024).
@@
IVs are commonly used in causal inference to address confounding. Recently, methods leveraging user search data as IVs have emerged, with the IV4Rec framework by Si et al. (Si et al. 2022) being a notable example. IV4Rec uses user search data to decompose user-item representations into causal and non-causal components, addressing some limitations of causal graph-based methods and reducing bias. However, identifying valid user search data as IVs remains challenging. Unlike these approaches, our work focuses on learning the representations of CIVs and their conditional sets, which are less restrictive than standard IVs.
The Proposed CIV4Rec Method
@@
In this section, we first introduce the problem definition, then explain the feasibility and rationality of our method by the causal graph, and then introduce the four main steps of our method. The overall workflow of our proposed CIV4Rec is visualised in Figure 3.
@@
Figure 3: The overall structure of CIV4Rec consists of four main steps, labelled 1, 2, 3, and 4 in the diagram. First, during feature encoding, CIV4Rec uses a backbone (MF or LightGCN) to encode the input data into a latent space representation. Second, the representations of CIV $Z_{t}$ and its conditional set $Z_{c}$ are learned from the user interaction data $X$. Third, the representation of CIV is used to decompose the treatment variable $W_{u,i}$, obtain the causal relationship representation $\widehat{W}_{u,i}$, and fuse it with the conditional set representation $Z_{c}$ to get the reconstructed treatment variable $W_{u,i}^{re}$. Finally, $W_{u,i}^{re}$ is used for the click prediction.
Problem Definition
@@
The Proposed Causal DAG
In this work, we proposed a causal DAG $\mathcal{G}$ as shown in Figure 2 to represent the causal relationships between measured and latent variables. Let $\mathcal{G}_{\underline{W}}$ be the manipulated graph, obtained by deleting all arrows emerging from nodes in $W$ within $\mathcal{G}$. In $\mathcal{G}_{\underline{W}}$, $Z_{t}$ and $W$ are d-connected when conditioned on $Z_{c}$ because of the existence of the edge $Z_{t}\rightarrow W$. However, $Z_{t}$ and $Y$ are d-separated by $Z_{c}$ since $Z_{t}$, $Z_{c}$ and $U_{c}$ form a collider at $W$, and $Z_{c}$ blocks the path $Z_{t}\leftarrow Z_{c}\rightarrow Y$. Furthermore, the effect of $Z_{t}$ on $Y$ is mediated solely by $W$ through the causal path $Z_{t}\rightarrow W\rightarrow Y$. Therefore, $Z_{t}$ is a CIV, and $Z_{c}$ is its corresponding conditional set. Note that $Z_{c}$ may contain information about latent factors $U_{c}$ due to the complex relationships in interactive data, as indicated by the dashed line between $U_{c}$ and $Z_{c}$ (Wu et al. 2022; Cheng et al. 2024). Many existing works have shown that confounding factors can affect recommendation outcomes, specifically $Z_{c}\rightarrow Y$ (Si et al. 2022, 2023a, 2023b). Additionally, since $Z_{t}$ and $Z_{c}$ are derived from the user interaction data $X$, they will affect $W$, i.e., $Z_{t}\rightarrow W$ and $Z_{c}\rightarrow W$ in $\mathcal{G}$.
Based on the proposed causal DAG, we introduce a data-driven CIV method for debiasing recommender systems called CIV4Rec. The advantages of CIV4Rec over existing debiasing methods are twofold: it does not require the specification of IVs through domain knowledge, and it uses CIV, which is more general than IV-based methods for addressing confounding bias caused by latent factors.
The Concepts of Treatment Variable, CIV and Its Conditional Set

### v2

# Interaction-Data-guided Conditional Instrumental Variables for Debiasing Recommender Systems

@@ Abstract
It is often challenging to identify a valid instrumental variable (IV), although the IV methods have been regarded as effective tools of addressing the confounding bias introduced by latent variables. To deal with this issue, an Interaction-Data-guided Conditional IV (IDCIV) debiasing method is proposed for Recommender Systems, called IDCIV-RS. The IDCIV-RS automatically generates the representations of valid CIVs and their corresponding conditioning sets directly from interaction data, significantly reducing the complexity of IV selection while effectively mitigating the confounding bias caused by latent variables in recommender systems. Specifically, the IDCIV-RS leverages a variational autoencoder (VAE) to learn both the CIV representations and their conditioning sets from interaction data, followed by the application of least squares to derive causal representations for click prediction. Extensive experiments on two real-world datasets, Movielens-10M and Douban-Movie, demonstrate that IDCIV-RS successfully learns the representations of valid CIVs, effectively reduces bias, and consequently improves recommendation accuracy.
footnotetext: Corresponding author
1 Introduction
@@
With the rapid development of the Internet, the amount of information has exploded, making it increasingly difficult for users to sift through vast amounts of data to find content that aligns with their preferences Luo et al. (2025); Gao et al. (2021); Zhang (2021). Recommender systems have emerged as a critical solution to this problem by analysing user behaviour data to deliver personalised recommendations, thereby enhancing user engagement and satisfaction Wang et al. (2020). Recommender systems have become an integral part of many digital platforms, finding extensive applications across various domains, including e-commerce Shoja and Tabrizi (2019), streaming media Gomez-Uribe and Hunt (2015), and social networks Liao et al. (2022), significantly improving information retrieval efficiency and user experience.
The performance of recommendation systems is often affected by latent confounders that are neither directly observable nor reflected in historical user-item interactions. For instance, factors such as social influence or peer preferences can significantly shape user behavior but are rarely captured in logged data. This misalignment may cause to recommendation systems misinterpret user intent, ultimately degrading recommendation quality. Effectively addressing latent confounders is thus vital for enhancing both the accuracy and robustness of recommender systems.
@@
Figure 1: Three causal graphs illustrate the assumptions underlying existing work. The variables are defined as follows: $U$: user preferences, II: displayed items, $Y$: user feedback or outcome variables, $C_{i}$: confounding factors affecting items, $C$: confounding factors, $Z$: instrumental variables, $W$: treatment variables (i.e., embeddings of user-item pairs). (a) The causal graph represents the traditional recommendation model; (b) the causal graph represents the debiasing method using causality; (c) the causal graph representation of the debiasing method using standard IV.
@@
User behaviour data is crucial for recommender systems in predicting user preferences. Existing models often assume this data is unbiased and accurately reflects user preferences, meaning the data accurately reflects user preferences Lan et al. (2024). Based on this assumption, many methods have been proposed, such as Matrix Factorisation (MF) Koren et al. (2009) and Neural Network-based Collaborative Filtering (NCF) He et al. (2017). These methods achieve the goal of predicting user preferences by fitting user behaviour data, as depicted in the causal Directed Acyclic Graph (DAG) Pearl (2009); Cheng et al. (2022) shown in Figure 1 (a), which illustrates the causal relationships between user preferences $\left(U\right)$ and user feedback $\left(Y\right)$, as well as between displayed items $\left(I\right)$ and $Y$. However, in the real-world, user behaviour is inevitably influenced by various unobservable confounding factors, such as item popularity (e.g., frequent recommendation of an item limits users’ choices) and user psychology (e.g., choosing to watch a movie for socialising) Chen et al. (2023). These factors introduce biases like popularity and conformity bias, leading models to learn false correlations and reducing their ability to accurately predict preferences.
Causal inference Zhang et al. (2024); Li et al. (2024) has been applied to reduce bias in recommender systems by designing causal DAGs to model data generation, identify biases, and guide model design Wang et al. (2020); Gao et al. (2024). For instance, Zhang et al. Zhang et al. (2021) introduced a causal graph to analyse item popularity’s impact and proposed the PDA training paradigm to correct popularity bias. Similarly, Zheng et al. Zheng et al. (2021) addressed conformity bias with the DICE model, disentangling user interest from conformity. However, these methods rely on the assumption that the real data follows the designed causal graphs, which may not hold in practice Cai et al. (2024). Complex real-world confounding makes such assumptions difficult to satisfy, potentially limiting the models’ effectiveness.
@@
Latent confounders, which are unobserved variables that simultaneously affect both the treatment (e.g., recommendation process) and the outcomes, present significant challenges for debiasing recommender systems. Instrumental Variables (IVs) are a common solution to this issue Caner and Hansen (2004); Pearl (2009), with a valid IV (e.g., $Z$ in Figure 1(c)) satisfying three key criteria Pearl (2009): (i) relevance to the treatment variable; (ii) an exclusive impact on the outcome through the treatment; and (iii) no shared confounders with the outcome. Several IV-based methods have been proposed for recommendation settings to reduce latent bias without relying on strict causal graph assumptions Si et al. (2022, 2023a, 2023b). For example, IV4Rec Si et al. (2022) leverages self-collected search data as an IV to effectively address latent confounding. However, verifying the latter two IV conditions from observational data alone is practically infeasible Brito and Pearl (2012); Cheng et al. (2023b), making the identification of valid IVs particularly challenging.
@@
Figure 2: A causal DAG illustrating our proposed IDCIV-RS method for discovering CIVs and their conditional sets from observational data. $W$, $Y$, $X$, and $U_{c}$ are the treatment, outcome, set of measured pretreatment variables, and latent confounders between $W$ and $Y$, respectively. $Z_{t}$ and $Z_{c}$ denote the representations of the CIVs and their conditional sets learned from $X$.
@@
In causal inference, researchers have used conditional IVs (CIVs) to tackle the limitations of standard IVs Pearl (2009); Brito and Pearl (2012); Cheng et al. (2024a, 2023a). CIVs offer more relaxed application conditions than standard IVs. Recently, Cheng et al. Cheng et al. (2023b) developed a CIV method (CIV.VAE) based on the variational autoencoder (VAE) Kingma and Welling (2013); Schölkopf et al. (2021); Schölkopf (2022) model, which generates CIVs and their conditional sets from data, significantly relaxing the constraints of standard IVs. However, this method is designed for tabular data, and no attempt has been made to adapt it to interactive data in recommender systems.
@@
To address the challenge of using CIV in interactive data, we first propose a causal DAG, as shown in Figure 2, to represent the causal relationships between observed and latent variables in interaction data within recommender systems. Building on this causal DAG, we develop an interaction-data-guided conditional IV (IDCIV) debiasing method for recommender systems, called IDCIV-RS. Specifically, IDCIV-RS uses the embeddings of user-item pairs (including users with the selected items, i.e., positive samples, and users with their pre-selected items, i.e., negative samples) as the treatment variable $W$, the user feedback as the outcome $Y$, and the user interaction data (only positive samples) as the pretreatment variable $X$. We assume that at least one CIV exists within $X$, capturing latent information such as user search behaviours that lead to interactions. The assumption is reasonable because in real-world scenarios, users are often influenced by certain external factors that cause their interactions to be not fully consistent with their true preferences. For example, users may interact with items because of certain external incentives (e.g., search recommendations, promotions, etc.) even though these items do not exactly match their true preferences. IDCIV-RS employs a VAE to generate the representations of the CIV and its conditional set from $X$, denoted as $Z_{t}$ and $Z_{c}$, as shown in Figure 2. The contribution of our work is summarised below:
@@
  • We develop an interaction-data-guided conditional IV (IDCIV) debiasing method for recommender systems, called IDCIV-RS, for learning representations of CIVs and its conditional sets under the proposed causal DAG. To the best of our knowledge, this is the first work to generate representations of CIVs and their conditional sets from interaction data for mitigating bias in recommender systems.
@@
  • Extensive experiments on two real-world datasets demonstrate that IDCIV-RS achieves optimal debiasing results and recommendation performance compared to state-of-the-art causal debiasing methods.
2 Related Work
@@
In this section, we review the recommender methods most closely related to our IDCIV-RS, including traditional recommender methods and causal recommender methods.
2.1 Traditional Recommender Methods
@@
Traditional recommender methods, primarily based on Collaborative Filtering (CF), typically assume that user behaviour data is unbiased Koren et al. (2021). The mainstream approach is model-based CF, which trains a model on user behaviour data to recommend items that align with user preferences. A classic method is MF Koren et al. (2009), which decomposes the user-item rating matrix to predict preferences. However, MF assumes that unselected items are incompatible with user preferences, ignoring cases where users may not have encountered those items. To address this, Rendle et al. Rendle et al. (2012) proposed Bayesian Personalized Ranking (BPR), which assumes users prefer selected items over unselected ones, enhancing preference inference in MF.
@@
With the rise of deep learning Luo et al. (2024); Zhang et al. (2022), He et al. He et al. (2017) proposed NCF, which uses multi-layer perceptrons (MLPs) to model non-linear user preferences. To capture richer behavioural signals, Wang et al. Wang et al. (2019) introduced Neural Graph Collaborative Filtering (NGCF), leveraging Graph Convolutional Networks (GCNs) to embed user-item interactions. He et al. He et al. (2020) further simplified this with LightGCN, improving both efficiency and accuracy. However, these methods often overlook popularity bias, which can be amplified during training and skew recommendations toward popular items.
2.2 Causal Recommender Methods
@@
To mitigate biases in recommender systems, researchers have increasingly adopted causal inference techniques. Early approaches used the Inverse Propensity Score (IPS) Wang et al. (2021); Schnabel et al. (2016); Bottou et al. (2013) to reduce bias by assigning an inverse propensity score (e.g., the inverse of item popularity) to user-item interactions during training, balancing the influence of popular and less popular items. However, IPS methods often suffer from high variance and instability. Building on IPS success, researchers have explored causal graph-based methods Zheng et al. (2021); He et al. (2023); Zhang et al. (2021) that model the generation mechanisms of user behaviour. These methods design specific models to address biases like popularity and conformity bias. Yet, the presence of unobserved confounders in real-world data limits the effectiveness of these causal graph assumptions Cai et al. (2024).
@@
IVs are commonly used in causal inference to address confounding. Recently, methods leveraging user search data as IVs have emerged, with the IV4Rec framework by Si et al. Si et al. (2022) being a notable example. IV4Rec uses user search data to decompose user-item representations into causal and non-causal components, addressing some limitations of causal graph-based methods and reducing bias. However, identifying valid user search data as IVs remains challenging. Unlike these approaches, our work focuses on learning the representations of CIVs and their conditional sets, which are less restrictive than standard IVs.
3 The Proposed IDCIV-RS Method
@@
In this section, we first introduce the problem definition, then explain the feasibility and rationality of our method by the causal graph, and then introduce the four main steps of our method. The overall workflow of our proposed IDCIV-RS is visualised in Figure 3.
@@
Figure 3: The overall structure of IDCIV-RS consists of four main steps, labelled 1, 2, 3, and 4 in the diagram. First, during feature encoding, IDCIV-RS uses a backbone (MF or LightGCN) to encode the input data into a latent space representation. Second, the CSEM module is used to learn the CIV $Z_{t}$ and its condition set $Z_{c}$ from the user-item interaction data $X$. Third, the representation of CIV is used to decompose the treatment variable $W_{u,i}$, obtain the causal relationship representation $\widehat{W}_{u,i}$, and fuse it with the conditional set representation $Z_{c}$ to get the reconstructed treatment variable $W_{u,i}^{re}$. Finally, $W_{u,i}^{re}$ is used for the click prediction.
3.1 Problem Definition
@@
3.2 The Proposed Causal DAG
In this work, we proposed a causal DAG $\mathcal{G}$ as shown in Figure 2 to represent the causal relationships between the measured and latent variables. Let $\mathcal{G}_{\underline{W}}$ be the manipulated graph, obtained by deleting all arrows emerging from nodes in $W$ within $\mathcal{G}$. In $\mathcal{G}_{\underline{W}}$, $Z_{t}$ and $W$ are d-connected when conditioned on $Z_{c}$ because of the existence of the edge $Z_{t}\rightarrow W$. However, $Z_{t}$ and $Y$ are d-separated by $Z_{c}$ since $Z_{t}$, $Z_{c}$ and $U_{c}$ form a collider at $W$, and $Z_{c}$ blocks the path $Z_{t}\leftarrow Z_{c}\rightarrow Y$. Furthermore, the effect of $Z_{t}$ on $Y$ is mediated solely by $W$ through the causal path $Z_{t}\rightarrow W\rightarrow Y$. Therefore, $Z_{t}$ is a CIV, and $Z_{c}$ is its corresponding conditional set. Note that $Z_{c}$ may contain information about latent factors $U_{c}$ due to the complex relationships in interactive data, as indicated by the dashed line between $U_{c}$ and $Z_{c}$ Wu et al. (2022); Cheng et al. (2024b). Many existing works have shown that confounding factors can affect recommendation outcomes, specifically $Z_{c}\rightarrow Y$ Si et al. (2022, 2023a, 2023b). Furthermore, since $Z_{t}$ and $Z_{c}$ are derived from the user interaction data $X$, they will affect $W$, i.e., $Z_{t}\rightarrow W$ and $Z_{c}\rightarrow W$ in $\mathcal{G}$.
Based on the proposed causal DAG, we present IDCIV-RS, an interaction-data-guided CIV debiasing method for recommender systems. IDCIV-RS offers two key advantages over existing approaches: it eliminates the need for domain-specific IV specification and leverages CIVs, which provide a more general framework for mitigating confounding bias from latent factors.
3.3 The Concepts of Treatment Variable, CIV and Its Conditional Set